OpenStudy (anonymous):

What is 62.5% of 1/5?

OpenStudy (anonymous):

0.125 :-)

OpenStudy (anonymous):

12.5% = 1/8 62.5 = 5/8 so 1/5 * 5/8 = 1/8

OpenStudy (anonymous):

12.5 /100 = .125

OpenStudy (anonymous):

Thanks!